# Blown up by an equilateral: Poncelet triangles about the incircle and their degeneracies

### v1

@@ 6. Ellipticity of X(k), k=1, 7, 8
Figure 23. Three Poncelet families for which the loci of $X_{k}$, $k=1,7,8$ are conics: (i) top left: confocal pair; bottom left: generic circular caustic; (iii) right: Chapple’s porism (bicentric $n=3$). Note that in (ii) and (iii), $X_{1}$ is stationary (center of caustic). Also note that in (iii), the loci of $X_{7}$ and $X_{8}$ are circles.
Figure 24. Three Poncelet families for which the loci of $X_{k}$, $k=1,7,8$ are non-conics: (i) top left: pair of homothetic ellipses; bottom left: pair of ‘dual’ ellipses; (iii) right: the MacBeath family.
@@ 7.2. Iso-X(2)
Direct from the location of the stationary barycenter and the fact that $|X_{1}X_{8}|=3|X_{1}X_{2}|$ [36, Incenter, eqn.(11)]. In fact, $|X_{1}X_{8}|=3(c\,b/(2a)-c\,b/(3a))=c\,b/(2a)=|C_{2}|$. ∎
@@
Since $X_{2}$ is stationary, the family conserves $|X_{1}X_{2}|$, which for any triangle with sidelengths $l_{i}$ is given by [36, Triangle Centroid, eqn.(11)]:
@@
The foci of the caustic are $X_{3}$ and $X_{4}$, therefore $X_{2}$ must be fixed, $|X_{3}X_{2}|=|X_{3}X_{4}|/3$ [36, Triangle Centroid, eqn. (13)]. Consider also that MacBeath family are the intouch triangles of bicentric polygons (Chapple’s porism). The $X_{2}$ of the intouch triangle is the Weill point $X_{354}$ of the reference [18], shown to be stationary over the bicentric family [21, Thm. 4.1]. ∎
@@ 7.3. Focal-X(4)
A triangle’s polar circle is centered on $X_{4}$ and has squared radius given by [36, Polar Circle]:
@@
The incenter-orthocenter squared distance is given by [36, Incenter, eqn.(6]:
@@
The MacBeath Poncelet family is inscribed in a circle and circumscribes the MacBeath inconic, with foci at the circumcenter and its isogonal conjugate, the orthocenter. Its center is the Euler center $X_{5}$, the midpoint of $X_{3}$ and $X_{4}$ [36, MacBeath inconic].
@@
The first relation is direct from [36, Circumcenter, eqn.(6)], noting that the (i) the circumradius $R$ is fixed and (ii) the caustic foci $X_{3},X_{4}$ are stationary. The second one is obtained via CAS-simplification. The third one is direct from [36, Circumradius, eqn.(5)]. ∎
@@
The first is direct from $|X_{3}-X_{4}|^{2}=9R^{2}-\sum{l_{i}^{2}}=2r_{pol}^{2}+R^{2}$ [36, Orthocenter, eqn.(14)], noting that the MacBeath circumradius $R$ is fixed and the caustic foci $X_{3}$ and $X_{4}$ are stationary. For the dual family, $X_{4}$ is stationary at the common center. Given the first relation it would therefore be sufficient to prove that $|X_{3}(t)|^{2}-R^{2}(t)$ is constant. The final expression is obtained via CAS simplification. ∎
@@ 7.4. Iso-X(7)
The Adams circle of a triangle is centered on $X_{1}$ and has radius $R_{A}$ given by [36, Adams Circle]:
@@ Appendix B Locus behaviors
Appendix C The Contact family
We now consider phenomena of the family of contact triangles of the incircle-circumscribed Poncelet analyzed above, i.e., with vertices at the tangents of each side with the incircle. Referring to Figure 29, the contact family is Ponceletian: it is circle-inscribed and envelops the polar image of tangents to $\mathcal{E}$ with respect to the incircle, also a conic [1].
@@ Appendix C The Contact family
Let $T_{o}$ be a triangle inscribed in the unit circle, and $\mathcal{E}_{c}$ an inconic with foci $f,g$, written as complex numbers. Referring to Figure 30:
@@
A Poncelet family of triangles inscribed in the unit circle and circumscribing an ellipse with foci $f,g\in\mathbb{C}$ contains an equilateral triangle iff $1/f+1/g$ is on the unit circle.
@@
to do First, assume that $T_{o}$ is an equilateral triangle in this family. Let $\zeta\in\mathbb{T}$ be one of the vertices of $T_{o}$ so that the other vertices are $e^{\frac{2\pi}{3}i}\zeta$ and $e^{\frac{4\pi}{3}i}\zeta$. Using the symmetric parameterization described in Theorem 1, let $\lambda_{o}\in\mathbb{T}$ be the parameter corresponding to triangle $T_{o}$. Plugging in the vertices in the parameterization, we get
@@
$z_{1}z_{2}+z_{2}z_{3}+z_{3}z_{1}=$ $fg+\lambda(\overline{f}+\overline{g})$
$z_{1}z_{2}z_{3}=$ $\lambda$
@@
A triangle’s Kiepert parabola is tangent to the three sides of a triangle. Its focus is $X_{110}$, always on the circumcircle, and its directrix is the Euler line $X_{2}X_{3}$ [36].
Still referring to Figure 30:
Proposition 35.
@@
Proof.
to do ∎
Figure 33. The construction of Propositions 37 and 38. $X_{110}$ (of $T_{o}=ABC$) is stationary over Poncelet if $\mathcal{O}$ lies on $\mathcal{L}_{35}$ (dashed red), the perpendicular bisector of $X_{3}$ and $X_{5}$. A vertex $A_{eq}$ of the equilateral contained in the family is shown.
@@
Proof.
to do ∎
Let $\mathcal{T}_{o}^{\prime}$ be the family of tangential triangles to $\mathcal{T}_{o}$, i.e., whose sides are tangent to the circumcircle at the vertices of $\mathcal{T}_{o}$ [36, Tangential Triangle]. Clearly, this family circumscribes a circle and will be inscribed in an ellipse $\mathcal{E}$, called above $\mathcal{T}$.

### v2

@@ 6. Ellipticity of X(k), k=1, 7, 8
Figure 23. Three Poncelet families for which the loci of $X_{k}$, $k=1,7,8$ are conics: (i) top left: confocal pair, live; (ii) bottom left: generic circular caustic, live; (iii) right: Chapple’s porism (bicentric $n=3$), live. Note that in (ii) and (iii), $X_{1}$ is stationary (center of caustic). Also note that in (iii), the loci of $X_{7}$ and $X_{8}$ are circles.
Figure 24. Three Poncelet families for which the loci of $X_{k}$, $k=1,7,8$ are non-conics: (i) top left: pair of homothetic ellipses, live; bottom left: pair of ‘dual’ ellipses, live; (iii) right: the MacBeath family, live.
@@ 7.2. Iso-X(2)
Direct from the location of the stationary barycenter and the fact that $|X_{1}X_{8}|=3|X_{1}X_{2}|$ [37, Incenter, eqn.(11)]. In fact, $|X_{1}X_{8}|=3(c\,b/(2a)-c\,b/(3a))=c\,b/(2a)=|C_{2}|$. ∎
@@
Since $X_{2}$ is stationary, the family conserves $|X_{1}X_{2}|$, which for any triangle with sidelengths $l_{i}$ is given by [37, Triangle Centroid, eqn.(11)]:
@@
The foci of the caustic are $X_{3}$ and $X_{4}$, therefore $X_{2}$ must be fixed, $|X_{3}X_{2}|=|X_{3}X_{4}|/3$ [37, Triangle Centroid, eqn. (13)]. Consider also that MacBeath family are the intouch triangles of bicentric polygons (Chapple’s porism). The $X_{2}$ of the intouch triangle is the Weill point $X_{354}$ of the reference [18], shown to be stationary over the bicentric family [21, Thm. 4.1]. ∎
@@ 7.3. Focal-X(4)
A triangle’s polar circle is centered on $X_{4}$ and has squared radius given by [37, Polar Circle]:
@@
The incenter-orthocenter squared distance is given by [37, Incenter, eqn.(6]:
@@
The MacBeath Poncelet family is inscribed in a circle and circumscribes the MacBeath inconic, with foci at the circumcenter and its isogonal conjugate, the orthocenter. Its center is the Euler center $X_{5}$, the midpoint of $X_{3}$ and $X_{4}$ [37, MacBeath inconic].
@@
The first relation is direct from [37, Circumcenter, eqn.(6)], noting that the (i) the circumradius $R$ is fixed and (ii) the caustic foci $X_{3},X_{4}$ are stationary. The second one is obtained via CAS-simplification. The third one is direct from [37, Circumradius, eqn.(5)]. ∎
@@
The first is direct from $|X_{3}-X_{4}|^{2}=9R^{2}-\sum{l_{i}^{2}}=2r_{pol}^{2}+R^{2}$ [37, Orthocenter, eqn.(14)], noting that the MacBeath circumradius $R$ is fixed and the caustic foci $X_{3}$ and $X_{4}$ are stationary. For the dual family, $X_{4}$ is stationary at the common center. Given the first relation it would therefore be sufficient to prove that $|X_{3}(t)|^{2}-R^{2}(t)$ is constant. The final expression is obtained via CAS simplification. ∎
@@ 7.4. Iso-X(7)
The Adams circle of a triangle is centered on $X_{1}$ and has radius $R_{A}$ given by [37, Adams Circle]:
@@ Appendix B Locus behaviors
Appendix C Fixed circumcircle family with equilateral
Referring to Figure 29, the family of contact triangles of a fixed incircle family is also Ponceletian: it is inscribed in a fixed circle and envelops the polar image of tangents to $\mathcal{E}$ with respect to the incircle, also a conic [1]. In particular, it is easy to see that the original family contains an equilateral triangle if and only if the family of contact triangles contains an equilateral. Thus we shall now study Poncelet families of unit circle-inscribed triangles and the many phenomena that arise when they contain an equilateral triangle.
@@ Appendix C Fixed circumcircle family with equilateral
Let $\mathcal{E}_{c}$ be an inconic with foci $f,g\in\mathbb{C}$, such that there is a Poncelet family of triangles inscribed in the unit circle $\mathbb{T}$ and circumscribing $\mathcal{E}_{c}$. Referring to Figure 30:
@@
A Poncelet family of triangles inscribed in the unit circle and circumscribing an ellipse with foci $f,g\in\mathbb{C}$ contains an equilateral triangle iff $1/f+1/g$ is on the unit circle $\mathbb{T}$.
@@
First, assume that $T_{o}$ is an equilateral triangle in this family. Let $\zeta\in\mathbb{T}$ be one of the vertices of $T_{o}$ such that the other vertices are $e^{\frac{2\pi}{3}i}\zeta$ and $e^{\frac{4\pi}{3}i}\zeta$. Using the symmetric parameterization described in Theorem 1, let $\lambda_{o}\in\mathbb{T}$ be the parameter corresponding to the triangle $T_{o}$. Plugging in the vertices in the parameterization, we get
@@
$\implies\lambda_{o}=-\frac{(f+g)}{\overline{f}\overline{g}}\implies\left|-{\frac{(f+g)}{\overline{f}\overline{g}}}\right|=|\lambda_{o}|=1$
$\implies|f+g|=|fg|\implies\left|\frac{1}{f}+\frac{1}{g}\right|=\left|\frac{f+g}{fg}\right|=1$
so $1/f+1/g$ is on the unit circle $\mathbb{T}$, as desired.
Conversely, suppose that $1/f+1/g$ is on the unit circle, that is, $|1/f+1/g|=1$. Rearranging, we have $|f+g|=|fg|$. Define $\lambda_{o}:=-(f+g)/(\overline{f}\overline{g})$. Then
$|\lambda_{o}|=\left|-\frac{(f+g)}{\overline{f}\overline{g}}\right|=\frac{|f+g|}{|\overline{f}\overline{g}}=\frac{|f+g|}{|fg|}=1$
By definition, $f+g+\lambda_{o}\overline{f}\overline{g}=0$. Moreover,
$fg+\lambda_{o}(\overline{f}+\overline{g})=\lambda_{o}\overline{\lambda_{o}}fg+\lambda_{o}(\overline{f}+\overline{g})=\lambda_{o}(\overline{\lambda_{o}}fg+\overline{f}+\overline{g})=$
$\lambda_{o}\overline{\left(f+g+\lambda_{o}\overline{f}\overline{g}\right)}=\lambda_{o}\overline{0}=0$
Hence, letting $z_{1},z_{2},z_{3}$ take the values of the 3 roots of $\zeta^{3}=\lambda_{o}$, we see that these $3$ roots form an equilateral triangle that satisfies the equations of the symmetric parameterization of Theorem 1. Thus, this equilateral triangle is part of the Poncelet family, as desired.
@@
The Kiepert parabola of a triangle is tangent to the three sides of a triangle. Its focus is $X_{110}$, always on the circumcircle, and its directrix is the Euler line $X_{2}X_{3}$ [37].
Still referring to Figure 30, it follows from Lemma 5:
Corollary 13.
@@
Figure 33. The construction of Propositions 36 and 37. $X_{110}$ (of $T_{o}=ABC$) is stationary over Poncelet if $\mathcal{O}$ lies on $\mathcal{L}_{35}$ (dashed red), the perpendicular bisector of $X_{3}$ and $X_{5}$. A vertex $A_{eq}$ of the equilateral contained in the family is shown.
@@
Let $\mathcal{T}_{o}^{\prime}$ be the family of tangential triangles to $\mathcal{T}_{o}$, i.e., whose sides are tangent to the circumcircle at the vertices of $\mathcal{T}_{o}$ [37, Tangential Triangle]. Clearly, this family circumscribes a circle and will be inscribed in an ellipse $\mathcal{E}$, called above $\mathcal{T}$.